# Qiuyu Wang2,    Hao Ouyang2,    Hao Chen1,    Chunhua Shen1,3,2

### v1

@@ B.3 Sampling and Evaluation
Appendix C More Analysis
C.1 Analysis on MATH500 and SVAMP datasets
@@
C.2 Temporal Semantic Entropy over Varying Generated Length
@@
C.3 Analysis on the Average Entropy in Blocks
@@
Appendix D More Experimental Results
D.1 Ablations on Scoring Rules for Combining TSE with Accuracy Reward
In Sec. 4.2, we combine TSE with the accuracy reward using the proper scoring rule (Gneiting & Raftery, 2007). The purpose of these scoring rules is to promote truthful confidence assessments: they attain their minimum value when the predicted confidence $c$ precisely mirrors the actual probability that the model’s output $o$ corresponds to the correct answer $o^{*}$.
Here, we conduct an ablation study on different scoring rules. Specifically, we consider the following four forms:
@@ D.1 Ablations on Scoring Rules for Combining TSE with Accuracy Reward
+RFT entropy 72.6 77.9 80.8 31.8 36.2 38.6 88.0 88.7 90.0 41.4 44.5 53.1
quadratic 69.8 78.6 81.8 30.0 35.0 40.8 88.3 88.7 88.3 45.3 40.6 54.3
logistic 70.1 78.5 81.7 30.6 37.2 39.0 89.0 87.7 89.3 42.2 41.1 52.7
spherical 70.2 79.5 81.7 30.8 34.8 41.2 88.7 89.3 89.0 53.1 43.0 55.1
D.2 Temporal Self-Consistency Voting After RFT
@@
D.3 Failure Case Analysis
Model Sudoku
Method / Seq Len 128 256 512
baseline 12.2 6.7 5.5
+ temporal voting 12.5 6.1 2.8
+ RFT accuracy reward (d1) 23.2 17.8 12.7
negative TSE reward (ours) 15.2 9.4 3.3
combining both (ours) 27.5 27.8 16.6
As discussed in Sec. 6, our method may rely on the model’s initial performance to achieve further improvements. To investigate this limitation, we take the challenging Sudoku dataset as a case study. As shown in Table S3, directly applying our proposed temporal self-consistency voting and temporal consistency reinforcement with the negative TSE reward alone results in a noticeable performance drop on the Sudoku dataset. For example, the original accuracy of the base model for the generation length of 512 is 5.5%, while the voting method achieves 2.8%, reflecting a decline of 2.7%.
D.4 Examples of time oscillation
@@
A notable example is seen in Example 3 below. Initially, the model correctly calculates the total cups, per-hour revenue, and the necessary per-hour cost deduction to derive the right profit per hour (step 54). However, during the reasoning process, it introduces erroneous masking regarding key calculation elements—the per-hour cost to subtract. Even though the answer remains correct at first (step 61), the flawed reasoning trajectory eventually results in miscalculations and an incorrect final response (step 62).
@@
Example 2 Problem: Shiela bought five cell phones for \$150 each for a 3-month installment. A 2% interest will be charged for each unit. How much will Shiela pay each month for 3 months[Uncaptioned image] Step 59: <<reasoning>> Shiela bought five cell phones for \$150 each, so the total cost is $5\times 150=\$750$. The interest for each unit is 2%, so the total interest is $750\times 0.02=\$15$. SinceIELa will pay this amount over 3 months, the monthly payment is \$750 + \$15 = $750\backslash 1.=255. <<$/reasoning>> <<answer>> 255 <</answer>> Step 60: <<reasoning>> Shiela bought five cell phones for \$150 each, so the total cost is $5\times 150=\$750$. The interest for each unit is 2%, so the total interest is $750\times 0.02=\$15$. SinceIELa will pay this amount over 3 months, the monthly payment is \$750 + \$15 = \$<<.>>750/3=>>250<<.>>750/3=>>250. <</reasoning>> <<answer>> 250 <</answer>>

### v2

@@ C.3 Sampling and Evaluation
Appendix D More Analysis
D.1 Analysis on MATH500 and SVAMP datasets
@@
D.2 Temporal Semantic Entropy over Varying Generated Length
@@
D.3 Analysis on the Average Entropy in Blocks
@@
Appendix E Limitations
E.1 Discussions on Limitations
While our temporal self-consistency voting and post-training approach demonstrates effectiveness in many scenarios, it exhibits significant limitations when applied to tasks where the model’s intermediate predictions are consistently inaccurate. As discussed in Sec. E.2, for the Sudoku dataset, the average correctness across all intermediate generation steps remains exceedingly low (below 5%), making it difficult to reliably vote for the correct answer. Similarly, RFT relies on the model already achieving reasonably good performance to produce meaningful reward signals (Prabhudesai et al., 2025; Agarwal et al., 2025), though combining TSE with the accuracy reward may alleviate this issue to some extent. This underscores that our approach depends on the model’s inherent ability to generate correct or near-correct answers in the sampling trajectory.
E.2 Failure Case Analysis
Model Sudoku
Method / Seq Len 128 256 512
baseline 12.2 6.7 5.5
+ temporal voting 12.5 6.1 2.8
+ RFT accuracy reward (d1) 23.2 17.8 12.7
negative TSE reward (ours) 15.2 9.4 3.3
combining both (ours) 27.5 27.8 16.6
As discussed in Sec. E.1, our method may rely on the model’s initial performance to achieve further improvements. To investigate this limitation, we take the challenging Sudoku dataset as a case study. As shown in Table S1, directly applying our proposed temporal self-consistency voting and temporal consistency reinforcement with the negative TSE reward alone results in a noticeable performance drop on the Sudoku dataset. For example, the original accuracy of the base model for the generation length of 512 is 5.5%, while the voting method achieves 2.8%, reflecting a decline of 2.7%.
Appendix F More Experimental Results
F.1 Training A Unified Model Across Multiple Tasks
To evaluate the generalization ability of our approach, we validate the proposed method in a multi-task setting. Specifically, we jointly train a single model on four datasets: GSM8K, MATH500, SVAMP, and Countdown. To ensure balanced training across tasks, we subsample each dataset so that the number of training examples is equal for all tasks.
Table S2 summarizes the results. Across all four benchmarks, our combined reward method consistently outperforms the supervised fine-tuning baselines. On GSM8K, we observe clear gains at shorter sequence lengths, for example, with LLaDA-1.5 on length 128, accuracy improves from 69.8 to 73.7, yielding +3.9 points. On SVAMP, similar short-sequence gains are observed, from 81.7 to 88.7, a +7.0 points increase. For the more challenging MATH500, the best improvement occurs at LLaDA-1.5, length 512, where performance increases from 35.4 to 41.6 (+6.2 points). Finally, on Countdown, our method achieves dramatic improvements, for instance, with LLaDA-1.5 at length 512, accuracy climbs from 20.7 to 50.4, a striking +29.7 points.
These results demonstrate that our method not only scales well across different reasoning domains but also enhances robustness under varying sequence lengths, confirming its effectiveness as a unified reward design for multi-task learning.
Table S2: Unified Model Performance Across Multiple Tasks. A single model is trained jointly on GSM8K, MATH500, SVAMP, and Countdown.
GSM8K MATH500 SVAMP Countdown
Method / Seq Len 128 256 512 128 256 512 128 256 512 128 256 512
+ RFT accuracy reward (d1) 71.3 77.3 80.7 28.0 33.6 39.2 84.0 84.0 85.0 25.8 22.3 37.1
negative TSE reward (ours) 71.5 77.2 81.2 29.0 33.8 39.0 86.3 84.7 87.0 36.7 25.4 46.1
combined reward (ours) 72.4 78.8 80.4 28.8 35.2 40.2 88.7 87.0 86.0 37.1 27.7 46.9
+ RFT accuracy reward (d1) 72.5 79.2 81.7 28.6 35.0 39.0 89.0 87.7 85.3 30.5 24.6 41.8
negative TSE reward (ours) 72.5 79.2 81.4 29.0 32.6 38.2 89.3 85.3 86.7 37.3 44.5 34.0
combined reward (ours) 73.7 79.5 82.5 28.2 35.0 41.6 88.3 90.0 88.0 37.9 29.7 50.4
F.2 Ablations on Scoring Rules for Combining TSE with Accuracy Reward
In Sec. 4.2, we combine TSE with the accuracy reward using the proper scoring rule (Gneiting & Raftery, 2007). The purpose of these scoring rules is to promote truthful confidence assessments: they attain their minimum value when the predicted confidence $c$ precisely mirrors the actual probability that the model’s output $o$ corresponds to the correct answer $o^{*}$. An exception is the spherical version we use, which consistently encourages both correctness and higher certainty, without penalizing overconfidence in incorrect predictions. Here, we conduct an ablation study on different scoring rules. Specifically, we consider the following four forms:
@@ F.2 Ablations on Scoring Rules for Combining TSE with Accuracy Reward
+RFT entropy 71.2 79.2 83.4 27.8 36.6 41.2 86.7 90.3 89.0 44.1 45.3 58.6
quadratic 73.4 79.1 83.2 29.2 33.2 42.4 86.0 86.7 90.0 43.4 45.3 57.4
logistic 71.2 79.2 83.1 30.2 33.8 41.0 88.3 89.0 91.0 48.8 46.9 61.3
spherical 73.2 80.5 84.0 29.6 35.4 41.4 86.3 90.3 89.0 44.5 46.9 63.3
F.3 Temporal Self-Consistency Voting After RFT
@@
F.4 Examples of time oscillation
@@
A notable case is seen in Example 3. Initially, the model correctly calculates the total cups, per-hour revenue, and the per-hour cost deduction to derive the right profit per hour (step 54). However, during reasoning, it introduces erroneous masking regarding key calculation elements—the per-hour cost to subtract. Even though the answer remains correct at first (step 61), the flawed reasoning trajectory eventually results in miscalculations and an incorrect final response (step 62).
Appendix G Change Log
Here we list the main modifications introduced in arXiv v2 compared to v1:
  • Bug Fix in LLaDA-1.5 Evaluation. We identified a bug in the previous experiments on LLaDA-1.5, where LoRA modules trained on LLaDA-1.5 were mistakenly evaluated by loading them into the LLaDA model. Interestingly, this setting revealed a certain degree of generalization, as the LoRA trained on LLaDA-1.5 still improved LLaDA’s performance. In the current version, we corrected this bug by properly evaluating LoRA with the LLaDA-1.5 backbone. As a result, both our method and the reproduced baseline d1 show improved performance (see Table 2 and Table S3).
  • Unified Multi-Task Training. We added new results for Training a Unified Model Across Multiple Tasks, as presented in Sec. F.1.
  • Text Revision. The manuscript was revised to improve readability, with the main content condensed to nine pages for clarity and precision.
@@ Appendix G Change Log
Example 2 Problem: Shiela bought five cell phones for \$150 each for a 3-month installment. A 2% interest will be charged for each unit. How much will Shiela pay each month for 3 months[Uncaptioned image] Step 59: <<reasoning>> Shiela bought five cell phones for \$150 each, so the total cost is $\mathdollar 5\times 150=\mathdollar 750$. The interest for each unit is 2%, so the total interest is $\mathdollar 750\times 0.02=\mathdollar 15$. SinceIELa will pay this amount over 3 months, the monthly payment is \$750 + \$15 = $750\backslash 1.=255. <<$/reasoning>> <<answer>> 255 <</answer>> Step 60: <<reasoning>> Shiela bought five cell phones for \$150 each, so the total cost is $\mathdollar 5\times 150=\mathdollar 750$. The interest for each unit is 2%, so the total interest is $\mathdollar 750\times 0.02=\mathdollar 15$. SinceIELa will pay this amount over 3 months, the monthly payment is \$750 + \$15 = \$<<.>>750/3=>>250<<.>>750/3=>>250. <</reasoning>> <<answer>> 250 <</answer>>